OpenStudy (anonymous):

find the slope of the line passing through the given points: (1/2, 3/5) and (1/4,1/3)

OpenStudy (anonymous):

slope formula m=(y1-y2)/(x1-x2) applying it gives us 16/15

OpenStudy (anonymous):

IT IS 15

OpenStudy (anonymous):

Solve the following for y:\[\frac{y-\frac{3}{5}}{x-\frac{1}{2}}\text{=}\frac{y-\frac{1}{3}}{x-\frac{1}{4}}\]\[y=\frac{16 x}{15}+\frac{1}{15} \]The slope is 16/15 The attached Mathematica statement shows that both points lie on the line.